OpenStudy (anonymous):

Prove the following Identity: 1/cosTheta-cosTheta/1+sinTheta=tanTheta

OpenStudy (anonymous):

\[\frac{ 1 }{ \cos \theta }-\frac{ \cos \theta }{ 1+\sin \theta }=\tan\]

OpenStudy (anonymous):

So far, I combined the LHS then multiplied the numerator and denominator by the inverse of the denominator. Unfortunately, I'm not sure if I'm doing it right.

OpenStudy (irishboy123):

i'd have tø to LHS and cø/(1+sø) to rhs. denominators on LHS then both cosø. thne cross multiply.

OpenStudy (anonymous):

I'm sorry, there are icons that are not showing up.

OpenStudy (irishboy123):

screen refresh. it's a problem a lot of the time.

OpenStudy (irishboy123):

WITH MAC SPECIAL CHARACTERS AND TYPO'S REMOVED: i'd have moved tan(theta) to LHS and cos(theta)/(1+sin(theta)) to rhs. denominators on LHS then both cos(theta). then cross multiply.

OpenStudy (anonymous):

Sorry, it took awhile for my screen to refresh.

OpenStudy (anonymous):

Sorry if it's dim, but this is what I've done so far.

OpenStudy (anonymous):

I can move tan theta to the other side and solve from there? I thought I would have to find the answer on the LHS to get to the RHS

OpenStudy (irishboy123):

what i meant!